OpenStudy (anonymous):

The value of y varies directly with x, and y = - 18 when x = 6. Find y when x = - 8.

OpenStudy (tkhunny):

"y varies directly with x" is code for "y = kx". Just write it down and you'll be well on your way.

OpenStudy (anonymous):

find the ratio of x to y. in this case -3x = y because -3*6 = -18. you know the formula now (its y=-3x) so just plug in -8 for x and you'll get y

OpenStudy (anonymous):

The fact that y varies directly with x means that, when x is multiply by a factor k, y is multiplied by the same factor k. So, since x is multiplied by -4/3, y is multiplied by -4/3, making the final value of y as follows: \[ y = -18 \cdot - \dfrac {4}{3} = \boxed {24}. \]